# Full QCD with milder topological freezing

### v1

@@ Abstract
We simulate $N_{f}=2+1$ QCD at the physical point combining open and periodic boundary conditions in a parallel tempering framework, following the original proposal by M. Hasenbusch for $2d$ $\mathrm{CP}^{N-1}$ models, which has been recently implemented and widely employed in $4d$ $\mathrm{SU}(N)$ pure Yang–Mills theories too. We show that using this algorithm it is possible to achieve an impressive reduction of the auto-correlation time of the topological charge in dynamical fermions simulations both at zero and finite temperature up to two orders of magnitude, allowing to avoid topology freezing down to lattice spacings as fine as $a\sim 0.02$ fm. Therefore, this implementation of the Parallel Tempering on Boundary Conditions algorithm has the potential to substantially push forward the investigation of the QCD vacuum properties by means of lattice simulations.
@@ 2.1 Lattice QCD action
$Z_{\rm LQCD}=\int[dU]e^{-S_{\mathrm{YM}}[U_{r}]}\det\left\{\mathcal{M}^{(\rm stag)}_{l}[U]\right\}^{\frac{1}{2}}\det\left\{\mathcal{M}^{(\rm stag)}_{s}[U]\right\}^{\frac{1}{4}},$ (1)
@@ 2.2 The PTBC algorithm with dynamical fermions
At finite lattice spacing, there is, strictly speaking, no proper notion of topological sectors. However, as the continuum limit is approached, the free-energy barriers between different pseudo-topological sectors grow, eventually diverging and properly restoring disconnected topological sectors. This is, in a few words, the physical origin of topological freezing DelDebbio:2004xh . Since these free energy barriers are developed essentially by the gauge action, we do not expect the algorithmic efficiency of the PTBC algorithm to improve if the boundary conditions of the links entering the fermion determinant are altered, hence our choice. [^1] Our choice is similar in spirit to one taken in the original PTBC pure-Yang–Mills $\mathrm{SU}(N)$ paper Bonanno:2020hht for what concerns the addition of the $\theta$-term in parallel tempering simulations. In that study, only the standard $\theta=0$ Wilson gauge action took part in the tempering of boundary conditions, while the imaginary-$\theta$ term did not. Moreover, as we will see in a moment, it will simplify the calculation of the swap probability.
@@
$K_{\mu}^{(r)}(x)=\begin{cases}c(r),\qquad&\mu=1,\,\,\,x_{1}=N_{s}-1,\,\,\,0\leq x_{0},x_{2},x_{3}<L_{d},\\ 1,\qquad&\text{elsewhere},\end{cases}$ (6)
@@ 3.1 Zero-temperature results
In order to put this comparison in more quantitative terms, we can estimate the auto-correlation time of $Q^{2}$ through a standard binned jack-knife analysis:
$\tau(Q^{2})=\frac{1}{2}\left(\frac{\mathrm{Err}_{\rm binned}(Q^{2})}{\mathrm{Err}_{\rm naive}(Q^{2})}\right)^{2},$ (14)
where $\mathrm{Err}_{\rm binned}$ and $\mathrm{Err}_{\rm naive}$ represent, respectively, the binned and the naive statistical errors on $\braket{Q^{2}}$. The obtained auto-correlation times are reported in Tab. 2, again expressed in both cases in units of the number of RHMC steps, meaning that for PTBC we multiplied it by $N_{r}$. As previously discussed, for the $a=0.048$ fm point we find auto-correlation times in agreement with each other. For the lattice spacing with $a=0.0397$ fm, instead, we find a net gain of about a factor of 2 in terms of $\tau(Q^{2})$ once the overhead introduced by the replicas is kept into account.
@@
For this simulation point, we also tried a larger defect size $l_{d}\sim 0.2$ fm with $P\simeq 20\%$. Also in this case we observe that the obtained result for the susceptibility is in excellent agreement with those obtained with the other parallel tempering setups. However, the auto-correlation time of $Q^{2}$, once the overhead introduced by the replicas is taken into account, turns out to be larger compared to the standard algorithm. This is not surprising, actually it is expected. As a matter of fact, it is clear that increasing the defect size will improve the decorrelation of the topological charge in the OBC replica, but at the same time will require a large number of replicas to achieve the same fixed average swap acceptance rate (empirically, $N_{r}\sim L_{d}^{3/2}$, see also Bonanno:2020hht ; Bonanno:2024nba ), thus excessively long defects are expected to lead to worse performances compared to the standard algorithm. An analogous behavior was observed for PTBC simulations in the pure-gauge theory Bonanno:2020hht .
@@
Table 2: Summary of the results obtained for $T\simeq 0$. The integrated auto-correlation time $\tau(Q^{2})$ was computed through a standard jack-knife binned analysis and is expressed for both algorithms in units of RHMC steps, meaning that that the one obtained with the PTBC algorithm was multiplied by the number of replicas $N_{r}$.
@@ 3.2 Finite-temperature results
$a$ [fm] $L_{d}$ $\chi^{1/4}$ [MeV] (PTBC) $\tau(Q^{2})$ (PTBC) $\chi^{1/4}$ [MeV] (Standard) $\tau(Q^{2})$ (Standard)
@@
Table 4: Summary of the results obtained for $T\simeq 570$ MeV. The results obtained for $a=0.0343$ fm and $0.0286$ fm with the standard multicanonic algorithm are taken from Ref. Athenodorou:2022aay . The integrated auto-correlation time $\tau(Q^{2})$ was computed through a standard jack-knife binned analysis and is expressed for both algorithms in units of RHMC steps, meaning that that the one obtained with the PTBC algorithm was multiplied by the number of replicas $N_{r}$.
Refer to captionRefer to caption
@@
Figure 5: Comparison of the multicanonical Monte Carlo evolutions of the cooled lattice topological charge $Q$ obtained with the PTBC and the standard algorithm for all lattice spacings explored at $T=570$ MeV. In all cases, the horizontal axis was reported in a common scale and expressed in numbers of RHMC steps, meaning that for PTBC we multiplied the Monte Carlo time by the number of replicas $N_{r}$. The parallel tempering run for the simulation point with $a=0.0343$ fm refers to the setup with $L_{d}=1$.
@@
We start our discussion by comparing the efficiency of the PTBC and the standard multicanonic algorithms. In Fig. 4, we plot an example of the multicanonic bias potential employed to enhance the probability of visiting suppressed topological sectors, while in Fig. 5 we show the Monte Carlo evolutions of the lattice topological charge $Q$ obtained with both algorithms, again compared by reporting both histories on a common horizontal axis (which implies that the Monte Carlo time of PTBC has been multiplied by the number of replicas $N_{r}$). In all cases we observe an improved number of topological fluctuations, especially for the finest lattice spacing explored. Again, in order to quantify the improvement obtained with PTBC, we computed the integated auto-correlation time, expressed in both cases using the same units (i.e., the number of RHMC steps). Results are reported in Tab. 4.
@@
For the coarsest lattice spacing explored with the PTBC algorithm, $a=0.0343$ fm, we tried two different defect lengths, namely $L_{d}=1$ and $2$. In the latter case, the PTBC auto-correlation time (i.e., multiplied by $N_{r}$) turned out to be larger by about a factor of 2 with respect to the standard algorithm one. In the former case instead, we observe substantial better performances compared to the standard algorithm, obtaining a net gain of about a factor of 3 in terms of $\tau(Q^{2})$. This thus suggests that defects smaller than $aL_{d}\sim 0.07$ fm should be employed in this case. Thus, for the other runs at finer lattice spacings, we simply adopted $L_{d}=2$ in all cases.
@@
For what concerns $a=0.0286$ fm, we observed a net gain of a factor of $2$ in terms of computational power. This has allowed us to greatly reduce the error on the gluonic determination for this lattice spacing compared to the result of Ref. Athenodorou:2022aay . For the finest lattice spacing explored the gain attained with PTBC algorithm is even more impressive. Indeed, while for parallel tempering we find an auto-correlation time which is of the same order of that observed for the coarser simulation points explored, using the standard multicanonic algorithm the Monte Carlo Markov chain does not appear to be thermalized yet even after thousands of RHMC steps, and thus only upper bounds for $\tau(Q^{2})$ and $\chi^{1/4}$ could be set for this run. Thus, these results point out that the PTBC algorithm shows an improved scaling of the auto-correlation time as a function of $1/a$ with respect to the standard algorithm also in the finite-temperature case, analogously to what has been deduced from our zero-temperature runs.
@@ Acknowledgements
STRONG-2020 project, under grant agreement No 824093. This work has also been partially supported by the project “Non-perturbative aspects of fundamental interactions, in the Standard Model and beyond” funded by MUR, Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN), Bando 2022, grant 2022TJFCYB (CUP I53D23001440006). Numerical calculations have been performed on the Leonardo machine at Cineca, based on the agreement between INFN and Cineca, under projects INF23_npqcd and INF24_npqcd.

### v2

@@ Abstract
We simulate $N_{f}=2+1$ QCD at the physical point combining open and periodic boundary conditions in a parallel tempering framework, following the original proposal by M. Hasenbusch for $2d$ $\mathrm{CP}^{N-1}$ models, which has been recently implemented and widely employed in $4d$ $\mathrm{SU}(N)$ pure Yang–Mills theories too. We show that using this algorithm it is possible to achieve a sizable reduction of the auto-correlation time of the topological charge in dynamical fermions simulations both at zero and finite temperature, allowing to avoid topology freezing down to lattice spacings as fine as $a\sim 0.02$ fm. Therefore, this implementation of the Parallel Tempering on Boundary Conditions algorithm has the potential to substantially push forward the investigation of the QCD vacuum properties by means of lattice simulations.
@@ 2.1 Lattice QCD action
$Z_{\rm LQCD}=\int[dU]e^{-S_{\mathrm{YM}}[U]}\det\left\{\mathcal{M}^{(\rm stag)}_{l}[U]\right\}^{\frac{1}{2}}\det\left\{\mathcal{M}^{(\rm stag)}_{s}[U]\right\}^{\frac{1}{4}},$ (1)
@@ 2.2 The PTBC algorithm with dynamical fermions
At finite lattice spacing, there is, strictly speaking, no proper notion of topological sectors. However, as the continuum limit is approached, the free-energy barriers between different pseudo-topological sectors grow, eventually diverging and properly restoring disconnected topological sectors. This is, in a few words, the physical origin of topological freezing DelDebbio:2004xh . Our expectation is that these free energy barriers are developed essentially by the gauge action, while the Dirac term should develop at most a step when jumping between two pseudo-topological sectors, associated to a non-zero eigenvalue becoming very small (due to the index theorem). For this reason, we do not expect the algorithmic efficiency of the PTBC algorithm to improve if the boundary conditions of the links entering the fermion determinant are altered, hence our choice. [^1] Our choice is similar in spirit to one taken in the original PTBC pure-Yang–Mills $\mathrm{SU}(N)$ paper Bonanno:2020hht for what concerns the addition of the $\theta$-term in parallel tempering simulations. In that study, only the standard $\theta=0$ Wilson gauge action took part in the tempering of boundary conditions, while the imaginary-$\theta$ term did not. Moreover, as we will see in a moment, it will simplify the calculation of the swap probability.
@@
$K_{\mu}^{(r)}(x)=\begin{cases}c(r),\qquad&\mu=1,\qquad x_{1}=N_{s}-1,\qquad 0\leq x_{0},x_{2},x_{3}<L_{d},\\ \\ 1,\qquad&\text{elsewhere},\end{cases}$ (6)
@@ 3.1 Zero-temperature results
In order to put this comparison in more quantitative terms, it is useful to compute the integrated auto-correlation time of $Q^{2}$. In this study, we will use the following estimator, which allows to evaluate the integated auto-correlation time through a standard binned jack-knife analysis Berg:2004fd ; DelDebbio:2004xh :
$\tau_{\mathrm{int}}(Q^{2})=\frac{1}{2}\left[\frac{\Delta_{\rm B}\left(Q^{2}\right)}{\Delta_{\rm 1}\left(Q^{2}\right)}\right]^{2}-\frac{1}{2},$ (14)
where $\Delta_{\rm B}(Q^{2})$ and $\Delta_{1}(Q^{2})$ represent, respectively, the binned and the naive (i.e., obtained neglecting auto-correlations) statistical errors on $\braket{Q^{2}}$. [^2] Note that the definition of the estimator of the integrated auto-correlation time given in Eq. (14) leads to $\tau_{\mathrm{int}}\geq 0$, with 0 the minimum value. Some authors adopt the equivalent definition without the $-1/2$, leading to $\tau_{\mathrm{int}}\geq 0.5$. At the practical level, the naive error $\Delta_{1}$ is obtained with a unitary bin size. The binned error $\Delta_{\rm B}$ instead exhibits a plateau after a certain bin size threshold $\mathrm{B}\gtrsim\overline{\mathrm{B}}$. As discussed in Ref. DelDebbio:2004xh , systematics affecting this estimator are negligible when $\tau_{\mathrm{int}}/\mathrm{B}\ll 1$, and when $\tau_{\mathrm{int}}$ and $\mathrm{B}$ are both much smaller than the sample size, which we verified was always the case in our PTBC simulations. Thus, the error we assigned to the integrated auto-correlation time was assessed from the small variations observed in the plateau exhibited by $\tau_{\mathrm{int}}(Q^{2})$ as a function of $\rm B$.
The obtained auto-correlation times are reported in Tab. 2, again expressed in both cases in units of the number of RHMC steps, meaning that for PTBC we multiplied the value of $\tau_{\mathrm{int}}$ obtained from Eq. (14) by $N_{r}$. As previously discussed, for the $a=0.048$ fm point we find integrated auto-correlation times in agreement with each other. For the lattice spacing with $a=0.0397$ fm, instead, we find a net gain of about a factor of 2 in terms of $\tau_{\mathrm{int}}(Q^{2})$ once the overhead introduced by the replicas is kept into account.
@@
For this simulation point, we also tried a larger defect size $l_{d}\sim 0.2$ fm with $P\simeq 20\%$. Also in this case we observe that the obtained result for the susceptibility is in excellent agreement with those obtained with the other parallel tempering setups. However, the integrated auto-correlation time of $Q^{2}$, once the overhead introduced by the replicas is taken into account, turns out to be larger compared to the standard algorithm. This is not surprising, actually it is expected. As a matter of fact, it is clear that increasing the defect size will improve the decorrelation of the topological charge in the OBC replica, but at the same time will require a large number of replicas to achieve the same fixed average swap acceptance rate (empirically, $N_{r}\sim L_{d}^{3/2}$, see also Bonanno:2020hht ; Bonanno:2024nba ), thus excessively long defects are expected to lead to worse performances compared to the standard algorithm. An analogous behavior was observed for PTBC simulations in the pure-gauge theory Bonanno:2020hht .
@@
Table 2: Summary of the results obtained for $T\simeq 0$. The integrated auto-correlation time $\tau_{\mathrm{int}}(Q^{2})$ was computed via Eq. (14) and is expressed for both algorithms in units of RHMC steps, meaning that the one obtained with the PTBC algorithm was multiplied by the number of replicas $N_{r}$.
@@ 3.2 Finite-temperature results
$a$ [fm] $L_{d}$ $\chi^{1/4}$ [MeV] (PTBC) $\tau_{\mathrm{int}}(Q^{2})$ (PTBC) $\chi^{1/4}$ [MeV] (Standard) $\tau_{\mathrm{int}}(Q^{2})$ (Standard)
@@
Table 4: Summary of the results obtained for $T\simeq 570$ MeV. The results obtained for $a=0.0343$ fm and $0.0286$ fm with the standard multicanonic algorithm are taken from Ref. Athenodorou:2022aay . The integrated auto-correlation time $\tau_{\mathrm{int}}(Q^{2})$ was computed via Eq. (14) and is expressed for both algorithms in units of RHMC steps, meaning that the one obtained with the PTBC algorithm was multiplied by the number of replicas $N_{r}$. The points out that, in that run with the standard algorithm, our estimate of $\tau_{\mathrm{int}}(Q^{2})$ is very rough, thus, the statistical error on $\chi$ is very likely to be underestimated.
@@
Figure 5: Top left plot: random walk of a configuration among the different replicas. We display how the boundary condition parameter $c(r)$ and the topological charge $Q$ change as a function of the Monte Carlo time, as well as the number of jumps of the topological charge as a function of $c(r)$. Top right and bottom plots: comparison of the multicanonical Monte Carlo evolutions of the cooled lattice topological charge $Q$ obtained with the PTBC and the standard algorithm for all lattice spacings explored at $T=570$ MeV. In all cases, the horizontal axis was reported in a common scale and expressed in numbers of RHMC steps, meaning that for PTBC we multiplied the Monte Carlo time by the number of replicas $N_{r}$. The parallel tempering run for the simulation point with $a=0.0343$ fm refers to the setup with $L_{d}=1$.
@@
We start our discussion by comparing the efficiency of the PTBC and the standard multicanonic algorithms. In Fig. 4, we plot an example of the multicanonic bias potential employed to enhance the probability of visiting suppressed topological sectors, while in Fig. 5 we show the Monte Carlo evolutions of the lattice topological charge $Q$ obtained with both algorithms, again compared by reporting both histories on a common horizontal axis (which implies that the Monte Carlo time of PTBC has been multiplied by the number of replicas $N_{r}$). From a visual inspection of the Monte Carlo histories, we observe that approaching the continuum limit parallel tempering outperforms the standard algorithm.
This is in particular clearly exemplified by the top left plot of Fig. 5, where we show how the topological charge $Q$ changes while a given configuration performs its random walk through the different replicas, experiencing different boundary conditions, parametrized by $c(r)$. As it can be seen, the topological charge is more likely to jump between two different sectors when $c(r)$ is closer to 0 (OBCs), and less likely when $c(r)$ is closer to 1 (PBCs). This means that the relevant source of topological fluctuations observed in the PTBC Monte Carlo histories is due to the opening of the boundary conditions, and not just a mere illusion introduced by the swaps among different replicas. To put this observation on a more quantitative ground, we computed the Pearson correlation coefficient $C=\left(\braket{xy}-\braket{x}\braket{y}\right)/\left(\sqrt{\braket{x^{2}}-\braket{x}^{2}}\sqrt{\braket{y^{2}}-\braket{y}^{2}}\right)$ between the boundary condition coefficient $x_{i}=c_{i}(r)$ and the jump of the topological charge $y_{i}=|\Delta Q_{i}|=|Q_{i+1}-Q_{i}|$, with $i$ the PTBC step, finding $C=-0.072(18)$. This result quantifies that it exists a non-vanishing correlation between the jumps of the topological charge and the opening of the boundaries on the defect region.
In order to better quantify the improvement obtained with PTBC with respect to the standard algorithm, we computed the integrated auto-correlation time of $Q^{2}$, expressed in both cases using the same units (i.e., the number of RHMC steps). Results are reported in Tab. 4. It is important to stress here that these integrated auto-correlation times refer to Eq.(14) calculated using the errors on $\braket{Q^{2}}$ obtained after reweighting. The motivation for this choice stems from the fact that we want a definition of $\tau_{\mathrm{int}}(Q^{2})$ which directly reflects the statistical accuracy on our target quantity (namely, the susceptibility after reweighting the topological bias potential away). [^3] For a related discussion on $\tau_{\mathrm{int}}(Q^{2})$ in the context of Jarzinsky-inspired reweighting in out-of-equilibrium simulations see Ref. Bonanno:2024udh .
@@
For the coarsest lattice spacing explored with the PTBC algorithm, $a=0.0343$ fm, we tried two different defect lengths, namely $L_{d}=1$ and $2$. In the latter case, the PTBC auto-correlation time (i.e., multiplied by $N_{r}$) turned out to be larger by about a factor of 2 with respect to the standard algorithm one. In the former case instead, we observe substantial better performances compared to the standard algorithm, obtaining a net gain of about a factor of 3 in terms of $\tau_{\mathrm{int}}(Q^{2})$. This thus suggests that defects smaller than $aL_{d}\sim 0.07$ fm should be employed in this case. Thus, for the other runs at finer lattice spacings, we simply adopted $L_{d}=2$ in all cases.
@@
For what concerns $a=0.0286$ fm, we observed a net gain of a factor of $2$ in terms of computational power. This has allowed us to greatly reduce the error on the gluonic determination for this lattice spacing compared to the result of Ref. Athenodorou:2022aay . For the finest lattice spacing explored the gain attained with PTBC algorithm is even more impressive. Indeed, while for parallel tempering we find an auto-correlation time which is of the same order of that observed for the coarser simulation points explored, using the standard multicanonic algorithm we are only able to make a rough estimate of $\tau_{\mathrm{int}}(Q^{2})$. Interestingly, we find that, with the standard algorithm, the integrated auto-correlation time turns out to be larger by at least a factor of 4 compared to PTBC. Thus, these results point out that the PTBC algorithm shows an improved scaling of the auto-correlation time as a function of $1/a$ with respect to the standard algorithm also in the finite-temperature case, analogously to what has been deduced from our zero-temperature runs.
@@ Acknowledgements
STRONG-2020 project, under grant agreement No 824093. G. C. acknowledges support from the National Centre on HPC, Big Data and Quantum Computing – SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU – National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N. 1.4 – CUP N. I53C22000690001. L. M. acknowledges support by the French Centre national de la recherche scientifique (CNRS) under an Emergence@INP 2023 project. This work has also been partially supported by the project “Non-perturbative aspects of fundamental interactions, in the Standard Model and beyond” funded by MUR, Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN), Bando 2022, grant 2022TJFCYB (CUP I53D23001440006). Numerical calculations have been performed on the Leonardo machine at Cineca, based on the agreement between INFN and Cineca, under projects INF23_npqcd and INF24_npqcd.